help please: Deanna is five more than six times older than Ted. Twice the sum of their ages is 94. How old is Ted?
Answer by Paul(988)

Ted's age = x
Deanna's age = five (5) more than (+) six (6) times Ted's age (6(x))
TWICE the sum of there ages = Ted's 2(x)
Deanna's = 2[5+6(x)]
Total ages = 94

2x+2(6x+5)=94
2x+12x+10=94
14x+10-10=94-10
14x=84
x=6 (Ted's age)
Denna's = 5+6(6) = 41

Check: 2(6+41)=94
94=94

Hence Ted's age is 6, and Deanna's age is 41.
Paul.